# Hilbert space methods for approximating multi-output latent variable Gaussian processes

### v2

@@ 3.3.2 Latent variable HSGPs
The vector of latent inputs $\mathbf{x}=(x_{i},\ldots,x_{n})$ is then passed to the approximation step in Eq.12 (and subsequently Eq.13 for the multi-output case). We call the resulting model (multi-output) latent variable HSGPs.
@@ 3.4 Bayesian inference
[…]
@@
The inference framework specified above is same for both exact and HSGPs. Posterior samples of $x$ and $\theta$ for all output dimensions are obtained via MCMC sampling, specifically adaptive Hamiltonian Monte Carlo (Neal, , 2011; Hoffman and Gelman, , 2014). We implemented the exact GP and the HSGP models in Stan using the RStan interface (Stan Development Team, , 2024).
@@ 3.5 Approximate latent variable GPs using VI
Approximating latent variable GPs using VI methods (VIGPs) usually include a combination of approximating both the covariance function and the posterior. Approximating the covariance function is carried out using inducing points to obtain a reduced rank representation of the Gram matrix (Rasmussen and Williams, , 2006). There are many inducing point methods (Quinonero-Candela and Rasmussen, , 2005), however, based on the suggestions of Bauer et al., (2016), we specifically consider the implementation discussed in Titsias, (2009). In this framework, the estimation of latent variables $x$ and GP hyperparameters $\theta$ after marginalizing over $f$ is carried out via the marginal likelihood
@@
Following the specifications discussed in Titsias and Lawrence, (2010), the joint probability distribution for the $d^{th}$ output dimension is given by
@@
where $M_{d}$ is the dimension-specific set of inducing points. The posterior $p(f_{d},M_{d}\mid y_{d})=p(f_{d}\mid M_{d},y_{d})\,p(M_{d}\mid y_{d})$ is subsequently approximated by a sparse variational distribution of the form
@@
where $q_{M}()$ is the variational distribution over the inducing points $M_{d}$. Model inference is carried out by the usual method of computing the variational lower bound (see Titsias and Lawrence, (2010) for further details). While the latent variable GPs were earlier designed for classification tasks (Lawrence, (2003), Lawrence, (2005)), VIGPs have later been used for GP regressions as well, thus making it an ideal candidate for reference and comparison to our proposed HSGPs. In this paper, we follow the implementations for VIGPs in Titsias and Lawrence, (2010); Ahmed et al., (2019) using the GP-LVM module of Pyro (Bingham et al., , 2019).
@@ 4 Simulation Study
For real-world data, we lack the ground truth values of latent variables. In contrast, in a simulated setting, we have full control over the data generating process including the true latent values. Thus, it is crucial to investigate the statistical properties of latent variable models using simulated data. To this end, we investigate exact GPs, HSGPs and VIGPs under various simulation scenarios exploring their behavior in terms of model convergence, model calibration, and estimation accuracy. The data generating conditions are inspired by Mukherjee et al., (2025).
@@ 4.1 Data generating scenarios
In all scenarios, we sample the between-dimension (output) correlation matrix $C\sim\text{LKJ}(\eta=1$) (Lewandowski et al., , 2009). This implies $C$ to be uniformly distributed within the set of all valid correlation matrices of dimension $D$. For the sample size $N$, we consider three cases $N=20,50\text{ }\text{and}\text{ }200$. We generate the ground truth latent inputs as $x_{i}\sim\text{Uniform}(0,10)$, where $i=1,\ldots,N$. Further, we assume a prior measurement SD of the noisy $\tilde{x}$ as $s=0.3$ (see Section 3.3). For the number of output dimensions $D$, we consider the three cases $D=5,10\text{ }\text{and}\text{ }20$. We perform 50 trials for each of the choices of $N$ and $D$ per simulation scenario.
@@ 4.2 Model specifications
We fit exact GPs, HSGPs, and VIGPs for all simulation scenarios, with the exception of omitting exact GPs for sample sizes higher than $N=20$ due to the scalability issues discussed earlier (see Section 1). For the HSGPs, we select the boundary conditions $L$ by multiplying a scaler adjustment $c=1.25$ to the range of the input prior $\tilde{x}$. This way, we ensure that we have a compact cover [L,L][-L,L] that prevents input values $x$ to be near the boundaries. By doing so, we avoid model convergence issues completely (see Section 4.3). We select the minimum number of basis functions $M_{\rm min}$ loosely based on the suggestions of Riutort-Mayol et al., (2022). The choices for $M_{\rm min}$ depends on the choice of covariance function to accurately approximate the degree of non-linearity. Thus, for the SE covariance function,
@@
Prior specifications for the exact GPs and HSGPs are kept same for all model parameters. Our prior choices are aligned with the data generating process since that is required for testing model calibration (see Section 4.4). Thus, we specify priors for length-scale $\rho\sim\text{Normal}^{+}(1,0.05^{2})$, GP marginal SD $\alpha\sim\text{Normal}^{+}(3,0.25^{2})$ and error SD $\sigma\sim\text{Normal}^{+}(1,0.25^{2})$. For the periodic data with higher oscillation, we specify a lower prior mean of $\rho\sim\text{Normal}^{+}(0.5,0.05^{2})$. Lastly, in case of the SE simulation scenario with highly varying length-scale, we specify a prior of $\rho\sim\text{Normal}^{+}(1,0.25^{2})$.
@@
Following the suggestions of the number of inducing points discussed in Titsias, (2009), we use $M=10$ inducing points for the VIGPs. In this framework, a small constant is added to the diagonal terms of the matrix. This helps maintain numerical stability while computing the GP covariance matrix. In case of VIGPs, we tested different values of diagonal constants starting from $\delta=10^{-12}$ (usually suggested default). Based on our experience and data generating conditions, the VIGPs need $\delta=10^{-3}$ for the GP covariance matrices to reliably become numerically positive definite. Higher number of inducing points would require an even larger $\delta$. Since $\delta=10^{-3}$ is already quite large, we keep $M=10$ to not further increase the bias caused by large values of $\delta$. As in the other methods, we set the prior measurement SD to $s=0.3$. The VIGP implementation in Pyro does not support priors on the GP hyperparameters, so we kept them at their uniform defaults.
@@
Both exact GPs and HSGPs are implemented in Stan (Stan Development Team, , 2024) and fitted with a single MCMC chain of 2000 iterations of which the first 1000 are discarded as warm-up. As shown in Mukherjee et al., (2025), model convergence is similar for multiple chains when fitting latent variable GPs. Thus, we run only a single chain per model to reduce overall computation times. The simulation studies are conducted on 50 vCPUs (Intel(R) Xeon(R) Gold 6230R CPU @ 2.10 GHz) with 720 GB work memory. The average runtime per dataset for the HSGPs with $N=200$ and $D=20$, the most computationally expensive case, was approximately 28 minutes for the SE model, 42 minutes for the Matern 3/2 model, and 34 minutes for the Matern 5/2 model. In the same scenarios, the VIGPs took approximately 2 minutes regardless of the covariance function. The exact GPs (just for $N=20$ and $D=20$) took 48 minutes for the SE model, 1.6 hours for the Matern 3/2 and 2.4 hours for the Matern 5/2 model. The HSGPs were significantly faster than the exact GPs but slower than VIGPs. The slower speed of HSGPs compared to VIGPs is counter-balanced by highly consistent model calibrations, superior estimation accuracy for latent variables, and overall much more stable model performance, as shown below.
@@ 4.3 Model convergence
We investigate MCMC convergence of our fitted exact GPs and HSGPs for all of the six simulation study scenarios discussed before. We use standard MCMC sampling diagnostics including state-of-the-art versions of the scale reduction factor $\widehat{R}$, bulk effective sample size (Bulk-ESS) and tail effective sample size (Tail-ESS) (Vehtari et al., , 2021). A combined check of these measures provide a comprehensive picture of the parameter-specific model convergence. In general, $\widehat{R}$ should be very close to 1 and should ideally not exceed 1.01 (Vehtari et al., , 2021). We additionally consider a more relaxed threshold of 1.1 in our simulation studies, since the ground truth is available as another layer of evaluation. Bulk-ESS indicates the reliability of measures of central tendency such as the posterior mean or median. Tail-ESS indicates the reliability of the 5% and 95% quantile estimates, which are then used to construct credible intervals. Both Bulk-ESS and Tail-ESS should have values greater than 100 times the number of MCMC chains (higher is better). We compute all the convergence measures with the posterior package (Bürkner et al., , 2023). Inference of the VIGPs is done via numerical optimization, which we considered to have converged if no convergence warnings were issued by the optimizer.
Figure 1: Squared exponential scenario: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs and HSGPs. The y-axes for Bulk and Tail ESS plots are log10 transformed.
In Figure 1, we show $\widehat{R}$, Bulk-ESS, and Tail-ESS for the latent $x$ and GP hyperparameters of both exact GPs and HSGPs with SE covariance function. While the exact GPs reach and exceed the relaxed $\widehat{R}$ threshold of 1.1 for some simulated datasets, the HSGPs consistently satisfy the much stricter 1.01 threshold of model convergence. HSGPs subsequently also have much higher Bulk and Tail-ESS. Overall, based on the diagnostics, HSGPs show much more consistent and stable convergence as compared to exact GPs. We provide convergence diagnostic figures for the other simulation scenarios in the supplementary materials Section C (Figures S3-S7)where we see similar results.
@@ 4.4 Testing model calibration
We use simulation based calibration (SBC) (Modrák et al., , 2023; Talts et al., , 2020) to test model calibration for estimating latent inputs $x$. Through SBC, we check if the model is able to produce posterior distributions that are consistent with the data generating process, thus correctly accounting for the implied uncertainty.
@@
To that end, briefly, SBC aims to test the goodness of posterior approximations by exploiting self-consistency properties of Bayesian models. This results in a rank statistic that is uniformly distributed under the assumption of a well calibrated model (Modrák et al., , 2023), which can subsequently be checked both graphically and numerically. As a graphical test, we plot the empirical cumulative distribution function (ECDF) of the ranks along with their 95% confidence regions under the assumption of uniformity (Säilynoja et al., , 2022). ECDFs lying outside of their confidence region indicate miscalibrations. These graphical tests also yield a test statistic (the log $\gamma$ score; see Säilynoja et al., , 2022), which we further analyze numerically. More details on SBC is provided in supplementary materials Section A.
@@
Figure 3: Squared exponential scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
In Figure 2, we show the rank ECDFs obtained from the exact GPs, HSGPs, and VIGPs for the SE simulation scenario ($N=20$ and $D=20$). The results indicate that the latent variable estimates obtained from the HSGPs are consistently well calibrated. In contrast, for exact GPs, we see miscalibrations for some latent variables shown by their corresponding rank sample ECDFs lying outside the threshold region. In case of VIGPs, the instances of miscalibrations are even more severe.
@@
To streamline calibration checking across simulation scenarios, we analyze the log $\gamma$ scores using a multi-level model (Bürkner, , 2017) described in supplementary materials Section A. In Figure 3, we show the predicted log $\gamma$ scores of the different GP models under the SE simulation scenario.
@@
Figure 4: Squared exponential scenario (highly varying $\rho$): log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
For all the choices of $N$ and $D$, the latent variable estimates from HSGPs are consistently well calibrated. By comparison, the exact GPs fail the calibration test for $D=20$. The VIGPs shows miscalibrations especially for higher output dimension cases throughout all the choices of $N$. Additionally, exact GPs exhibit worse calibrations (decreasing log $\gamma$ scores) when estimating larger values of the latent inputs. This issue directly causes problems in estimating latent values closer to the boundary. VIGPs show similar behavior for lower sample sizes $N=20$ although not as strong as exact GPs. The HSGPs are able to overcome and rectify this behavior. For the special scenario of SE data with highly variable length-scales as shown in Figure 4, VIGPs and HSGPs with small number of basis function struggle to pass the calibration tests, especially for $N=200$. The miscalibrations are however alleviated with a higher number of basis functions in HSGPs. The log $\gamma$ results for the Matern 3/2 and 5/2 simulation scenarios are presented in supplementary materials Section D (Figures S8 and S9).

### v3

@@ 3.3.2 Latent variable HSGPs
The vector of latent inputs $\mathbf{x}=(x_{i},\ldots,x_{N})$ is then passed to the approximation step in Eq.(12) (and subsequently Eq.(13) for the multi-output case). We call the resulting model (multi-output) latent variable HSGPs.
@@ 3.4 Bayesian inference
where $p(y_{d}\mid x,\theta_{d})$ denotes the GP-based likelihood for a single output dimension and p(xx~)p($\mid\tilde{x}$) denotes the prior for the latent $x$ implied by the measurement model in Eq.(16). The details of prior specifications used in our experiments are further discussed in Section 4. Using Bayes’ rule, we obtain the joint posterior over $x$ and $\theta$ as
@@
The inference framework specified above is same for both exact and HSGPs. Posterior samples of $x$ and $\theta$ for all output dimensions are obtained via MCMC sampling, specifically adaptive Hamiltonian Monte Carlo (Neal, 2011; Hoffman and Gelman, 2014). We implemented the exact GP and the HSGP models in Stan using the RStan interface (Stan Development Team, 2024).
@@ 3.5 Approximate latent variable GPs using VI
Approximating latent variable GPs using VI methods (VIGPs) usually include a combination of approximating both the covariance function and the posterior. Approximating the covariance function is carried out using inducing points to obtain a reduced rank representation of the Gram matrix (Rasmussen and Williams, 2006). There are many inducing point methods (Quinonero-Candela and Rasmussen, 2005), however, based on the suggestions of Bauer et al. (2016), we specifically consider the implementation discussed in Titsias (2009). In this framework, the estimation of latent variables $x$ and GP hyperparameters $\theta$ after marginalizing over $f$ is carried out via the marginal likelihood
@@
Following the specifications discussed in Titsias and Lawrence (2010), the joint probability distribution for the $d^{th}$ output dimension is given by
@@
where $M_{d}$ is the dimension-specific set of inducing points. The inducing points are designated as variational parameters and optimized through variational inference Titsias (2009); Titsias and Lawrence (2010). The posterior $p(f_{d},M_{d}\mid y_{d})=p(f_{d}\mid M_{d},y_{d})\,p(M_{d}\mid y_{d})$ is subsequently approximated by a sparse variational distribution of the form
@@
where $q_{M}()$ is the variational distribution over the inducing points $M_{d}$. Model inference is carried out by the usual method of computing the variational lower bound (see Titsias and Lawrence (2010) for further details).
While the latent variable GPs were earlier designed for classification tasks (Lawrence (2003), Lawrence (2005)), VIGPs have later been used for GP regressions as well, thus making it an ideal candidate for reference and comparison to our proposed HSGPs. In this paper, we follow the implementations for VIGPs in Titsias and Lawrence (2010); Ahmed et al. (2019) using the GP-LVM module of Pyro (Bingham et al., 2019) where the user only needs to set the number of inducing points and the internal working takes care of their optimization for sparse latent input estimation.
@@ 4 Simulation Study
For real-world data, we lack the ground truth values of latent variables. In contrast, in a simulated setting, we have full control over the data generating process including the true latent values. Thus, it is crucial to investigate the statistical properties of latent variable models using simulated data. To this end, we investigate exact GPs, HSGPs and VIGPs under various simulation scenarios exploring their behavior in terms of model convergence, model calibration, and estimation accuracy. The data generating conditions are inspired by Mukherjee et al. (2025).
@@ 4.1 Data generating scenarios
In all scenarios, we sample the between-dimension (output) correlation matrix $C\sim\text{LKJ}(\eta=1$) (Lewandowski et al., 2009). This implies $C$ to be uniformly distributed within the set of all valid correlation matrices of dimension $D$. For the sample size $N$, we consider three cases $N=20,50\text{ }\text{and}\text{ }200$. Due to the computational complexity of exact GPs, we only consider them for $N=20$ case. Further, a special case for the GP data scenarios is also demonstrated for the HSGPs and VIGP with $N=1000$ in Section 4.6. Unlike the smaller $N$ cases, here the fitted models aren’t aligned with the true data generating process, that is, we specify vague priors for our HSGP hyperparameters for $N=1000$ case. We generate the ground truth latent inputs as $x_{i}\sim\text{Uniform}(0,10)$, where $i=1,\ldots,N$. Further, we assume a prior measurement SD of the noisy $\tilde{x}$ as $s=0.3$ (see Section 3.3). For the number of output dimensions $D$, we consider the three cases $D=5,10\text{ }\text{and}\text{ }20$. We perform 50 trials for each of the choices of $N$ and $D$ per simulation scenario.
@@ 4.2 Model specifications
We fit exact GPs, HSGPs, and VIGPs for all simulation scenarios, with the exception of omitting exact GPs for sample sizes higher than $N=20$ due to the scalability issues discussed earlier (see Section 1). For the HSGPs, we select the boundary conditions $L$ such that it contains the input space completely. This is obtained by multiplying a scalar adjustment $c=1.25$ (Riutort-Mayol et al., 2022) to the range of the input prior $\tilde{x}$. Thus, we ensure that we have a compact cover [L,L][-L,L] that prevents input values $x$ to be near the boundaries. By doing so, we ensure stable MCMC model convergence for the latent variable inputs (see Section 4.3). We select the minimum number of basis functions $M_{\rm min}$ loosely based on the suggestions of Riutort-Mayol et al. (2022). The choices for $M_{\rm min}$ depends on the choice of covariance function to accurately approximate the degree of non-linearity. Thus, for the SE covariance function,
@@
Prior specifications for the exact GPs and HSGPs are kept same for all model parameters. Our prior choices are aligned with the data generating process since that is required for testing model calibration (see Section 4.4). Thus, we specify priors for length-scale $\rho\sim\text{Normal}^{+}(1,0.05^{2})$, GP marginal SD $\alpha\sim\text{Normal}^{+}(3,0.25^{2})$ and error SD $\sigma\sim\text{Normal}^{+}(1,0.25^{2})$. For the periodic data with higher oscillation, we specify a lower prior mean of $\rho\sim\text{Normal}^{+}(0.5,0.05^{2})$. Lastly, in case of the SE simulation scenario with highly varying length-scale, we specify a prior of $\rho\sim\text{Normal}^{+}(1,0.25^{2})$. The number of estimated primary parameters of interest for each HSGPs are provided in Table 1.
Table 1: Primary parameters of interest
Name Notation Number of parameters
Latent inputs $x$ NN
Length scale $\rho$ $D$
GP marginal SD $\alpha$ $D$
Error SD $\sigma$ $D$
  • Note: $N$ stands for number of observations or sample size. $D$ denotes the number of output dimensions.
@@
Following the suggestions of the number of inducing points discussed in Titsias (2009), we use $M=10$ inducing points for the VIGPs. In this framework, a small constant is added to the diagonal terms of the matrix. This helps maintain numerical stability while computing the GP covariance matrix. In case of VIGPs, we tested different values of diagonal constants starting from $\delta=10^{-12}$ (usually suggested default). Based on our tests, for 10 inducing variables (points), the VIGPs need at least $\delta=10^{-3}$ for a numerically positive definite GP covariance matrices and reliably finish an entire simulation scenario. Higher number of inducing points would require an even larger $\delta$. Since $\delta=10^{-3}$ is already quite large, we keep $M=10$ to not further increase the bias caused by large values of $\delta$. As in the other methods, we set the latent prior measurement SD to $s=0.3$. As for the GP hyperparameters, since the VIGP implementation in Pyro does not support custom priors we keep them at their uniform defaults for the simulation scenarios with $N=20,50\text{ }\text{and}\text{ }200$. For the $N=1000$ (see Section 4.6), we introduce a VIGP case where GP hyperparameters are fixed to the data generating conditions, essentially creating a favorable condition for VIGPs.
@@
Both exact GPs and HSGPs are implemented in Stan (Stan Development Team, 2024) and fitted with a single MCMC chain of 2000 iterations of which the first 1000 are discarded as warm-up. As shown in Mukherjee et al. (2025), model convergence is similar for multiple chains when fitting latent variable GPs. Thus, we run only a single chain per model to reduce overall computation times. The simulation studies are conducted on 50 vCPUs (Intel(R) Xeon(R) Gold 6230R CPU @ 2.10 GHz) with 720 GB work memory. The average runtime per dataset for the HSGPs with $N=200$ and $D=20$, the most computationally expensive case, was approximately 28 minutes for the SE model, 42 minutes for the Matern 3/2 model, and 34 minutes for the Matern 5/2 model. In the same scenarios, the VIGPs took approximately 2 minutes regardless of the covariance function. The exact GPs (just for $N=20$ and $D=20$) took 48 minutes for the SE model, 1.6 hours for the Matern 3/2 and 2.4 hours for the Matern 5/2 model. The HSGPs were significantly faster than the exact GPs but slower than VIGPs. The slower speed of HSGPs compared to VIGPs is counter-balanced by highly consistent model calibrations, superior estimation accuracy for latent variables, and overall much more stable model performance, as shown below.
@@ 4.3 Model convergence
We investigate MCMC convergence of our fitted exact GPs and HSGPs for all of the six simulation study scenarios discussed before. We use standard MCMC sampling diagnostics including state-of-the-art versions of the scale reduction factor $\widehat{R}$, bulk effective sample size (Bulk-ESS) and tail effective sample size (Tail-ESS) (Vehtari et al., 2021). A combined check of these measures provide a comprehensive picture of the parameter-specific model convergence. In general, $\widehat{R}$ should be very close to 1 and should ideally not exceed 1.01 (Vehtari et al., 2021). We additionally consider a more relaxed threshold of 1.1 in our simulation studies, since the ground truth is available as another layer of evaluation. Bulk-ESS indicates the reliability of measures of central tendency such as the posterior mean or median. Tail-ESS indicates the reliability of the 5% and 95% quantile estimates, which are then used to construct credible intervals. Both Bulk-ESS and Tail-ESS should have values greater than 100 times the number of MCMC chains (higher is better). We compute all the convergence measures with the posterior package (Bürkner et al., 2023). Inference of the VIGPs is done via numerical optimization, which we considered to have converged if no convergence warnings were issued by the optimizer.
Figure 1: Squared exponential scenario: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs (for $N=20$) and HSGPs (combined for $N=20,50\text{ }\text{and}\text{ }200$ cases). The y-axes for Bulk and Tail ESS plots are log10 transformed.
In Figure 1, we show $\widehat{R}$, Bulk-ESS, and Tail-ESS for the latent $x$ and GP hyperparameters of both exact GPs and HSGPs with SE covariance function. While the exact GPs reach and exceed the relaxed $\widehat{R}$ threshold of 1.1 for some simulated datasets, the HSGPs consistently satisfy the much stricter 1.01 threshold of model convergence. HSGPs subsequently also have much higher Bulk and Tail-ESS. Overall, based on the diagnostics, HSGPs show much more consistent and stable convergence as compared to exact GPs. We provide convergence diagnostic figures for the other simulation scenarios in the supplementary materials Section C (Figures S3-S7) where we see similar results. We additionally provide further analysis for MCMC convergence of latent inputs and GP hyperparameters compared between exact GPs and HSGPs in supplementary materials Section D (Figures S8-S11).
@@ 4.4 Testing model calibration
We use simulation based calibration (SBC) (Modrák et al., 2023; Talts et al., 2020) to test model calibration for estimating latent inputs $x$. Through SBC, we check if the model is able to produce posterior distributions that are consistent with the data generating process, thus correctly accounting for the implied uncertainty.
@@
To that end, briefly, SBC aims to test the goodness of posterior approximations by exploiting self-consistency properties of Bayesian models. This results in a rank statistic that is uniformly distributed under the assumption of a well calibrated model (Modrák et al., 2023), which can subsequently be checked both graphically and numerically. As a graphical test, we plot the empirical cumulative distribution function (ECDF) of the ranks along with their 95% confidence regions under the assumption of uniformity (Säilynoja et al., 2022). ECDFs lying outside of their confidence region indicate miscalibrations. These graphical tests also yield a test statistic (the log $\gamma$ score; see Säilynoja et al., 2022), which we further analyze numerically. More details on SBC is provided in supplementary materials Section A.
@@
Figure 3: Squared exponential scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity, that is, models with values less than 0 are miscalibrated. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
In Figure 2, we show the rank ECDFs obtained from the exact GPs, HSGPs, and VIGPs for the SE simulation scenario ($N=20$ and $D=20$). The results indicate that the latent variable estimates obtained from the HSGPs are consistently well calibrated. In contrast, for exact GPs, we see miscalibrations for some latent variables shown by their corresponding rank sample ECDFs lying outside the threshold region. In case of VIGPs, the instances of miscalibrations are even more severe. The rank ECDFs for Matern 3/2 and 5/2 scenarios are reported in supplementary materials Section E (Figures S12 and S13) where the results are same as the SE scenario.
@@
To streamline calibration checking across simulation scenarios , we analyze the log $\gamma$ scores using a multi-level model (Bürkner, 2017) described in supplementary materials Section A. In Figure 3, we show the predicted log $\gamma$ scores of the different GP models under the SE simulation scenario. This is equivalent to summarizing the different cases of rank ECDFs across all simulation scenarios.
@@
Figure 4: Squared exponential scenario (highly varying $\rho$): log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity , that is, models with values less than 0 are miscalibrated. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
For all the choices of $N$ and $D$, the latent variable estimates from HSGPs are consistently well calibrated. By comparison, the exact GPs fail the calibration test for $D=20$ as their log $\gamma$ scores remain consistently less than the critical value corresponding to the 95% confidence threshold. This can be primarily attributed to the computational issues of exact GPs for latent variable inputs as seen through MCMC convergence diagnostics in Section 4.3 along with supplementary materials Section C and D. The VIGPs shows miscalibrations especially for higher output dimension cases throughout all the choices of $N$. Additionally, exact GPs exhibit worse calibrations (decreasing log $\gamma$ scores) when estimating larger values of the latent inputs. This issue directly causes problems in estimating latent values closer to the boundary. VIGPs show similar behavior for lower sample sizes $N=20$ although not as strong as exact GPs. The HSGPs are able to overcome and rectify this behavior. For the special scenario of SE data with highly variable length-scales as shown in Figure 4, VIGPs and HSGPs with small number of basis function struggle to pass the calibration tests, especially for $N=200$. The miscalibrations are however alleviated with a higher number of basis functions in HSGPs. The log $\gamma$ results for the Matern 3/2 and 5/2 simulation scenarios are presented in supplementary materials Section E (Figures S14 and S15).